# Ultrasonic Medical Tissue Imaging Using Probabilistic Inversion: Leveraging Variational Inference for Speed Reconstruction and Uncertainty Quantification

### v2

@@ 3.1 Overview of SVGD
$\bm{m}_{i}^{t+1}=\bm{m}_{i}^{t}+\epsilon^{t}\left[k\left(\bm{m}_{i}^{t},\bm{m}_{j}^{t}\right)\nabla_{\bm{m}_{i}^{t}}\log p\left(\bm{m}_{i}^{t}\mid\bm{d}^{\mathrm{obs}}\right)+\nabla_{\bm{m}_{i}^{t}}k\left(\bm{m}_{i}^{t},\bm{m}_{j}^{t}\right)\right].$ (24)
The above iterative update process, using the kernel function and the gradient of the log posterior, refines the approximation to the true posterior distribution.
As demonstrated in Ref. [34], SVGD effectively employs particles to approximate the mean and standard deviation of complex distributions, including Gaussian Mixture Model (GMM). GMMs are widely used to represent complex probability distributions due to their flexibility in capturing multimodal behaviors. The GMM distribution $p(x)$ is defined as a weighted sum of Gaussian components,
$p(x)=\sum_{i}w_{i}p_{i}(x)=\sum_{i}w_{i}\mathcal{N}(x;\mu_{i},\sigma_{i}),$ (25)
[…]
$p_{i}(x)=\frac{1}{\sqrt{2\pi\sigma_{i}^{2}}}\exp\left(-\frac{(x-\mu_{i})^{2}}{2\sigma_{i}^{2}}\right).$ (26)
The mean of the overall GMM is calculated as
$\mu_{\text{GMM}}=\sum_{i}w_{i}\mu_{i},$ (27)
which reflects the weighted contribution of each component’s mean. The standard deviation of the GMM is
$\sigma_{\text{GMM}}=\sqrt{\sum_{i}w_{i}\left(\sigma_{i}^{2}+\mu_{i}^{2}\right)-\mu_{\text{GMM}}^{2}}.$ (28)
The formulation (28) captures the combined spread of the distribution, accounting for both the variances of individual components and the deviation of their means.
Figure 1: SVGD approximates the mean and standard deviation of the Gaussian Mixture Model (GMM). (a) The black line represents the density distribution of the GMM, with mean $\mu_{\text{GMM}}=0.9$ and standard deviation $\sigma_{\text{GMM}}=1.446$. The blue, orange, and green lines correspond to the three components of the GMM, with weights $w_{1}=0.2$, $w_{2}=0.5$, and $w_{3}=0.3$, respectively. (b) The initial distribution of the particles, generated from the standard normal distribution $\mathcal{N}(x;0,1)$. (c) The results of the SVGD algorithm. The dashed red line shows the ground truth mean. The mean and standard deviation of the particles are $0.883$ and $1.314$, respectively. (d) The MSE of the mean and standard deviation plotted against the iterations of the SVGD algorithm.
Mathematically, any continuous probability distribution can be approximated arbitrarily well by a GMM given a sufficient number of components [40]. However, for complex target distributions, optimizing the means, variances, and weights of the Gaussian components becomes computationally expensive as the number of components increases. SVGD offers a nonparametric approach to approximate distributions by iteratively moving particles to better align with the target distribution. SVGD is advantageous for scenarios where the target distribution is intractable, as it eliminates the need for explicit parameter optimization. Consequently, if SVGD effectively approximate a GMM, it could also be leveraged to approximate other complex distributions with better efficiency. In the following, an example is employed to illustrate the effectiveness of SVGD in approximating the mean and standard deviation of a GMM.
Assuming a GMM with three components, the model is defined as
$p(x)=0.2\mathcal{N}(x;-1,1)+0.5\mathcal{N}(x;1,0.5^{2})+0.3\mathcal{N}(x;2,1.5^{2}),$ (29)
and the corresponding density and its components are illustrated in Figure 1 (a). By applying Eqs. (27) and (28), the mean and standard deviation of the GMM are
$\mu_{\text{GMM}}=0.9,\ \sigma_{\text{GMM}}=1.446,$ (30)
which represent the ground truth values for the SVGD algorithm to approximate. The SVGD algorithm begins with $n$ particles generated from the standard Gaussian distribution $\mathcal{N}(x;0,1)$, as depicted in Figure 1 (b). The initial random distribution of the particles shows noticeable differences from the GMM distribution shown in Figure 1 (a). To apply the SVGD algorithm to the GMM, the approximated mean and standard deviation are denoted as $\hat{u}$ and $\hat{\sigma}$, respectively, and the fit is evaluated using the mean squared error (MSE) between the true and approximated values. Following the SVGD update steps in Eq. (24), the final distribution of the particles is shown in Figure 1 (c), where most particles are centered around the ground truth mean. The computed mean and standard deviation of the particles are $\hat{u}=0.883$ and $\hat{\sigma}=1.314$, respectively, which are close to the ground truth values. Figure 1 (d) displays the MSE of the mean over iterations, with the inset showing the variations in the MSE of the standard deviation. After 600 iterations, both the MSE of the mean and standard deviation converge to their optimal values.
3.2 Gradient Calculation of Log-Posterior for SVGD Method
@@
To evaluate the gradient of $\log p(\bm{m})$, it is assumed that the prior distribution $p(\bm{m})$ follows a Gaussian distribution with mean vector $\bm{\mu}$ and covariance matrix $\mathbf{\Sigma}$,
$p(\bm{m})=\frac{1}{(2\pi\det\mathbf{\Sigma})^{\frac{d}{2}}}\exp\left(-\frac{1}{2}(\bm{m}-\bm{\mu})^{\top}\mathbf{\Sigma}^{-1}(\bm{m}-\bm{\mu})\right),$ (37)
where $\mathbf{\Sigma}^{-1}$ is the inverse of the covariance matrix. To simplify the calculation, the mean-field Gaussian parameterisation is employed [33]. The prior distribution $p(\bm{m})$ is approximated by a factorized Gaussian distribution over the parameters $\bm{m}=(m_{1},m_{2},\dots,m_{d})$,
$p(\bm{m})=\prod_{i=1}^{d}p(m_{i}).$ (38)
In the above factorized form, the probability $p(m_{i})$ for each dimension is a Gaussian distribution
$p(m_{i})=\frac{1}{\sqrt{2\pi\sigma_{i}^{2}}}\exp\left(-\frac{(m_{i}-\mu_{i})^{2}}{2\sigma_{i}^{2}}\right),$ (39)
where $\mu_{i}$ and $\sigma_{i}^{2}$ are the mean and variance of the $i$-th component $m_{i}$. The factorization (38) reduces the complexity of the prior distribution by assuming that the parameters are independent. The gradient of $\log p(\bm{m})$ with respect to each $m_{i}$ is
$\nabla_{m_{i}}\log p\left(\bm{m}\right)=$ $\frac{\partial}{\partial m_{i}}\left(-\frac{1}{2}\log(2\pi\sigma_{i}^{2})-\frac{(m_{i}-\mu_{i})^{2}}{2\sigma_{i}^{2}}\right)=-\frac{m_{i}-\mu_{i}}{\sigma_{i}^{2}}.$ (40)
@@
4 Simulated Results
Figure 2: Inversion setup and initial conditions for the synthetic data experiment using a linear array transducer. (a) The true model setup. The orange dots and green diamonds indicate the locations of the sources and receivers, respectively. (b) The Ricker wavelet used as the source waveform. (c) The initial model for the conventional FWI. (d) An example of the initial particles for SVGD.
In this section, the gradients in Eqs. (36) and (40) are combined to apply the SVGD update rule (24) for solving FWI problems under in-silico experimental conditions. The primary focus is on the SOS distribution of objects or tissues within the acoustic field. Using the obtained SOS information, imaging of the objects or tissues is performed. To assess the convergence of the SVGD algorithm, Eq. (4) is used as the loss function. Once the update steps converge, the SOS distribution is visualized by plotting the mean of the particles, and the uncertainty in the corresponding SOS is measured through the standard deviation. Additionally, to evaluate the effectiveness of SVGD, the conventional FWI method is applied to obtain a deterministic result. In conventional FWI, the gradient is computed using the adjoint method, and the loss function (4) is minimized via gradient descent.
4.1 Linear Array Transducer
@@
First, the focus is on using synthetic data from a linear array transducer for inversion imaging. A two-dimensional grid model is employed, with a grid size of $201\times 201$ and a spatial step size of $0.5$ mm, resulting in a total computational area of $10\,\text{cm}\times 10\,\text{cm}$. As shown in Figure 2(a), the true model setup consists of a circular object located at the center of the computational area. The object is a homogeneous isotropic medium with a diameter of $1.5$ cm and an SOS of $1.7$ km/s. The background SOS is set at $1.5$ km/s, resembling the speed of sound under water. This difference in SOS results in variations in the wave propagation, providing a basis for the inversion algorithm to distinguish the target object from the background medium.
@@
The simulated experiment utilizes 21 transmitting sources and 101 receivers. In Figure 2(a), the sources are evenly distributed at $x=0.15$ cm, and the receivers are uniformly placed at $x=9.9$ cm. The starting and ending positions of both sources and receivers are $0.5$ cm away from the upper and lower boundaries of the computational area, respectively. The source waveform is set as the Ricker wavelet, with a central frequency of $0.5$ MHz and a duration time of $0.2$ ms. As illustrated in Figure 2(b), the frequency spectrum of the Ricker wavelet is well-concentrated, providing high energy within a narrow frequency band.
For the conventional FWI, the initial model sets the SOS within the grid to $1.5$ km/s, as shown in Figure 2(c). For the SVGD algorithm, Gaussian white noise with a mean of 00 and a standard deviation of $0.01$ km/s is added to the initial FWI model to generate different initial particles. Figure 2(d) displays one example of the initial particles. These initial models serve as starting points, which are iteratively updated during the inversion process. The prior distribution is assumed to follow the Gaussian distribution Eq. (37), with $\mu_{i}=1.5$ km/s and $\sigma_{i}=1$ km/s. During the inversion iterations, the Devito software (https://github.com/devitocodes/devito) is used to obtain forward data and the gradient from the adjoint-state method.
@@
Figure 3 presents the results of the inversion process. The SOS distribution obtained from the conventional FWI is shown in Figure 3(a). The circular object in the true model is reconstructed as an ellipse in the inversion result. This transformation is primarily due to insufficient angular coverage. In the case of the linear array transducer, the sources and receivers do not provide adequate angular coverage, limiting the acquisition of wave propagation data from multiple angles. As a result, the inversion algorithm is unable to accurately reconstruct the true shape of the target object. A similar transformation is observed in Figure 3(b), which shows the mean SOS distribution of the SVGD particles. Compared to the conventional FWI results, the mean SOS in SVGD particles is closer to the true model. The standard deviation, shown in Figure 3(c), measures the uncertainty in the mean SOS distribution. The central region of Figure 3(c) exhibits low uncertainty, indicating that the SVGD algorithm provides a more reliable SOS estimate in this area. Thus, the standard deviation serves to evaluate the accuracy and reliability of the mean result across different regions. Figure 3(d) plots the loss values against the iteration number, showing the convergence of the inversion process.
4.2 Ring Array Transducer
@@
Figure 4: Results of the inversion process for the ring array transducer simulation. (a) The true model setup. The orange dots and green diamonds indicate the locations of the sources and receivers, respectively. (b) The SOS distribution obtained from the conventional FWI. (c) The mean SOS from SVGD. (d) The standard deviation of the SOS distribution from SVGD. (e) The loss values plotted against the iteration number. (f) The relative error between the predicted mean SOS and the true SOS. The gray dashed circle (radius = 3.25 cm) indicates the region used to evaluate the effectiveness of SVGD.
To alleviate the effects of insufficient angular coverage caused by the linear array transducer setup, an alternative transducer configuration commonly used in medical ultrasound imaging is the ring array. The inversion imaging process is then performed using simulated data from the ring array transducer configuration. As shown in Figure 4(a), the true model consists of a rectangular object positioned at the center of the computational domain, with both width and length measuring $3$ cm. The central region has an SOS of $1.7$ km/s, with a $1.5$ cm × $1.5$ cm rectangular core, while the surrounding region has an SOS of $1.6$ km/s. The background SOS is set at $1.5$ km/s. The experiment uses $32$ sources and $32$ receivers, with all sources uniformly distributed on a circle with a diameter of 8.5 cm, centered at the grid’s center. The layout of the receivers is identical to that of the sources. The remaining parameters are the same as those in the linear array configuration shown in Figure 2.
[…]
5 Breast Data Results

### v3

@@ 2.3 SVGD-based posterior approximation and uncertainty quantification in FWI
$\bm{m}_{j}^{t+1}=\bm{m}_{j}^{t}+\frac{\epsilon^{t}}{n}\sum_{i=1}^{n}\left[k\left(\bm{m}_{i}^{t},\bm{m}_{j}^{t}\right)\nabla_{\bm{m}_{i}^{t}}\log p\left(\bm{m}_{i}^{t}\mid\bm{d}^{\mathrm{obs}}\right)+\nabla_{\bm{m}_{i}^{t}}k\left(\bm{m}_{i}^{t},\bm{m}_{j}^{t}\right)\right].$ (15)
Where $n$ is the number of particles. The above iterative update process, using the kernel function and the gradient of the log posterior, refines the approximation to the true posterior distribution.
@@
To evaluate the gradient of $\log p(\bm{m})$, the prior distribution $p(\bm{m})$ is defined as a truncated Gaussian with bounded support $m_{\min},m_{\max}]$, ensuring that $\bm{m}$ remains within its physically plausible range. Following the mean-field parameterization strategy [34, 44], the prior is expressed as
$p(\bm{m})=\prod_{i=1}^{d}p(m_{i}),\quad m_{i}\in[m_{\min},m_{\max}].$ (22)
Each marginal $p(m_{i})$ is specified as a truncated Gaussian,
$p(m_{i})\propto\frac{1}{\sqrt{2\pi}\sigma_{i}}\exp\!\left(-\frac{(m_{i}-\mu_{i})^{2}}{2\sigma_{i}^{2}}\right),$ (23)
normalized by the cumulative probability within the truncation interval. The mean and variance of the $i$-th component $m_{i}$ are denoted by $\mu_{i}$ and $\sigma_{i}^{2}$, respectively. The factorization (22) reduces the complexity of the prior distribution by assuming independence across parameters. The gradient of $\log p(\bm{m})$ with respect to each $m_{i}$ is
$\nabla_{m_{i}}\log p(\bm{m})=-\frac{m_{i}-\mu_{i}}{\sigma_{i}^{2}},\quad m_{i}\in[m_{\min},m_{\max}].$ (24)
@@
3 Results
Figure 1: Inversion setup and initial conditions for the synthetic data experiment using a linear array transducer. (a) The true model setup. The orange dots and green diamonds indicate the locations of the sources and receivers, respectively. (b) The Ricker wavelet used as the source waveform. (c) The initial model for the conventional FWI. (d) An example of the initial particles for SVGD, with a total of 20 particles used in the inversion.
In this section, the gradients in Eqs. (21) and (24) are combined to apply the SVGD update rule (15) for solving FWI problems under different experimental conditions. The primary focus is on the SOS distribution of objects or tissues within the acoustic field. Using the obtained SOS information, imaging of the objects or tissues is performed. Once the update steps converge, the SOS distribution is visualized by plotting the mean of the particles, and the uncertainty in the corresponding SOS is measured through the standard deviation. Additionally, to evaluate the effectiveness of SVGD, the conventional FWI method is applied to obtain a deterministic result. In conventional FWI, the gradient is computed using the adjoint method, and the loss function (4) is minimized via gradient descent.
3.1 Synthetic data imaging with linear array transducer
@@
First, the focus is on using synthetic data from a linear array transducer for inversion imaging. A two-dimensional grid model is employed, with a grid size of $301\times 301$ and a spatial step size of $\Delta x=\Delta y=0.3$ mm, resulting in a total computational area of $9\,\text{cm}\times 9\,\text{cm}$. As shown in Figure 1(a), the true model setup consists of a circular object located at the center of the computational area. The object is a homogeneous isotropic medium with a diameter of $1.5$ cm and an SOS of $1.7$ km/s. The background SOS is set at $1.5$ km/s, resembling the speed of sound under water. This difference in SOS results in variations in the wave propagation, providing a basis for the inversion algorithm to distinguish the target object from the background medium.
@@
The simulated experiment utilizes 31 transmitting sources and 101 receivers. In Figure 1(a), the sources are evenly distributed at $x=0.15$ cm, and the receivers are uniformly placed at $x=8.9$ cm. The starting and ending positions of both sources and receivers are $0.3$ cm away from the upper and lower boundaries of the computational area, respectively. The source signal is a Ricker wavelet with a central frequency of 0.5 MHz and a duration of 0.12 ms. As illustrated in Figure 1(b), the frequency spectrum of the Ricker wavelet is well-concentrated, delivering high energy within a narrow band. The time step $\Delta t$ is determined based on the Courant-Friedrichs-Lewy (CFL) stability condition, $\Delta t\leq C\Delta x/c_{\text{max}}$, where $c_{\text{max}}$ is the maximum speed of sound in the model and $C$ is a dimensionless constant [13]. With $C=0.46$, the time step is set to $\Delta t=8\times 10^{-5}$ ms, resulting in 1501 temporal samples. To ensure numerical accuracy and suppress dispersion, the spatial discretization satisfies a minimum of 10 grid points per wavelength.
For the conventional FWI, the initial model sets the SOS within the grid to $1.5$ km/s, as shown in Figure 1(c), resulting in an initial squared slowness of $\bm{m}=1/\bm{c}^{2}\approx 0.444~\text{s}^{2}/\text{km}^{2}$. The true SOS ranges from $c_{\min}=1.5$ km/s to $c_{\max}=1.7$ km/s, the squared slowness is constrained within $m_{\min}=1/c_{\max}^{2}$ and $m_{\max}=1/c_{\min}^{2}$ during the inversion. Values of $\bm{m}$ exceeding these bounds are truncated to prevent numerical issues such as NaN (not a number) during computation. The inversion is performed using the steepest descent optimization method with a maximum of 80 iterations. The step size is empirically scheduled to decrease with iterations, initially set to $9\times 10^{-3}$, and subsequently reduced by a decay factor of 0.57 every 20 iterations. The stopping criterion is satisfied when either the relative change in the objective function between consecutive iterations falls below $10^{-5}$ or the iteration count reaches 80.
For the SVGD algorithm, 20 initial particles are generated by perturbing the initial model of the conventional FWI with Gaussian white noise (mean 0, standard deviation 0.01). Figure 1(d) displays one example of the initial particles. The initial particles are iteratively updated according to the SVGD update rule Eq.(15). A truncated Gaussian prior Eq.(23) is adopted for the squared slowness, with $\mu_{i}=0.444~\text{s}^{2}/\text{km}^{2}$ and $\sigma_{i}=0.1~\text{s}^{2}/\text{km}^{2}$. The optimization procedure also employs the steepest descent method, using a maximum of 80 iterations. The step size is initialized at $9\times 10^{-3}$ and decays by a factor of 0.57 every 20 iterations. The inversion terminates when the relative change in the objective function between iterations drops below $10^{-5}$ or when the iteration limit is reached.
@@
Figure 2 presents the results of the inversion process, where the reconstructed SOS distribution is obtained via $\bm{c}=1/\sqrt{\bm{m}}$. The result of the conventional FWI is shown in Figure 2(a). The circular object in the true model is reconstructed as an ellipse in the inversion result. This transformation is primarily due to insufficient angular coverage. In the case of the linear array transducer, the sources and receivers do not provide adequate angular coverage, limiting the acquisition of wave propagation data from multiple angles. As a result, the inversion algorithm is unable to accurately reconstruct the true shape of the target object. A similar transformation is observed in Figure 2(b), which shows the mean SOS distribution of the SVGD particles. Compared to the conventional FWI results, the mean SOS in SVGD particles is closer to the true model. The standard deviation, shown in Figure 2(c), measures the uncertainty in the mean SOS distribution. The central region of Figure 2(c) exhibits low uncertainty, indicating that the SVGD algorithm provides a more reliable SOS estimate in this area. Thus, the standard deviation serves to evaluate the accuracy and reliability of the mean result across different regions. Figure 2(d) plots the loss values against the iteration number, showing the convergence of the inversion process.
3.2 Synthetic data imaging with ring array transducer
@@
Figure 3: Results of the inversion process for the ring array transducer simulation. (a) The true model setup. The orange dots and green diamonds indicate the locations of the sources and receivers, respectively. (b) The SOS distribution obtained from the conventional FWI. (c) The mean SOS estimated by SVGD using 20 particles. (d) The standard deviation of the SOS distribution from SVGD. The pink dashed circle (radius = 3.27 cm) indicates the ROI used to evaluate the effectiveness of the inversion algorithms.
To mitigate the limitations of insufficient angular coverage inherent in linear array transducer configurations, the experiment adopts a ring array layout commonly employed in medical ultrasound imaging. A two-dimensional grid model comprising $335\times 335$ points with a spatial resolution of $\Delta x=\Delta y=0.3$ mm defines a computational domain measuring $10.02\,\text{cm}\times 10.02\,\text{cm}$. As depicted in Figure 3(a), the true model includes a central rectangular object with dimensions of 3 cm × 3 cm. The innermost core, measuring 1.5 cm × 1.5 cm, has a SOS of 1.7 km/s, surrounded by a region with an SOS of 1.6 km/s, embedded in a background with an SOS of 1.5 km/s. A total of 32 sources and 32 receivers are evenly distributed along a circular ring with a diameter of 8.52 cm, centered within the grid. The receiver layout mirrors that of the sources. The excitation signal is a Ricker wavelet with a central frequency of 0.5 MHz and a duration of 0.12 ms. To ensure numerical stability, the time step $\Delta t$ adheres to the CFL condition with $C=0.3$, resulting in $\Delta t=8\times 10^{-5}$ ms and a total of 1601 temporal samples. The spatial discretization satisfies the requirement of at least 10 grid points per wavelength to minimize numerical dispersion.
For the conventional FWI, the initial model sets the SOS uniformly to $1.5$ km/s (the approximate SOS of pure water), resulting in an initial squared slowness of $\bm{m}=1/\bm{c}^{2}\approx 0.444~\text{s}^{2}/\text{km}^{2}$. The true SOS varies within the range $c_{\min}=1.5$ km/s to $c_{\max}=1.7$ km/s, and the squared slowness is accordingly constrained within $m_{\min}=1/c_{\max}^{2}$ and $m_{\max}=1/c_{\min}^{2}$ during the inversion. Any values of $\bm{m}$ that fall outside this range are truncated to ensure physical plausibility and to prevent numerical issues such as NaN during computation. The inversion is performed using the quasi-Newton optimization method L-BFGS-B, implemented via the scipy.optimize.minimize function in Python, with a maximum of 50 iterations. The algorithm employs an internal line search based on strong Wolfe conditions, with the step length adaptively chosen. The stopping criterion is met when the relative change in the objective function falls below $10^{-5}$ or when the number of iterations reaches the specified maximum.
For the SVGD algorithm, 20 initial particles are generated by adding Gaussian white noise (mean 0, standard deviation 0.01) to the initial FWI model. The particles are iteratively updated according to Eq. (15). The prior distribution is modeled as a truncated Gaussian distribution Eq.(23), with $\mu_{i}=0.444~\text{s}^{2}/\text{km}^{2}$ and $\sigma_{i}=0.1~\text{s}^{2}/\text{km}^{2}$. The optimization is carried out using the quasi-Newton method L-BFGS-B, implemented via the scipy.optimize.minimize function, with a maximum of 50 iterations. The algorithm incorporates an internal line search, and the step size is adaptively selected. The inversion terminates when the relative change in the objective function falls below $10^{-5}$ or when the iteration count reaches the maximum limit.
Figure 4: Loss and relative error analysis for the ring array transducer simulation. (a) The loss values plotted against the iteration number. (b) and (c) show the relative error for FWI and SVGD, respectively. The pink dashed circle (radius = 3.27 cm) indicates the ROI.
The reconstructed SOS is obtained by $\bm{c}=1/\sqrt{\bm{m}}$. For the SVGD method, $n$ particles ($n=20$) produce $n$ distinct reconstructions $\bm{c}_{i},~i=1,2,\cdots,n$, from which the mean and standard deviation of the SOS distribution are computed. Figure 3(b)-(c) show the SOS distribution obtained from the conventional FWI and SVGD, respectively. Both methods accurately reconstruct the true model, capturing the internal structure and contours of the object. Figure 3(d) measures the imaging uncertainty of the SVGD mean SOS. High uncertainty appears mainly in the green region of Figure 3(c). This area shows subtle SOS variations around 1.6 km/s, which also appear in the conventional FWI result Figure 3(b).
Figure 4(a) shows the loss values of FWI and SVGD as a function of iteration number, with SVGD converging more quickly. To quantify reconstruction accuracy, the relative error between the predicted model $\hat{\bm{m}}$ and the true model $\bm{m}$, defined as $(\hat{\bm{m}}-\bm{m})/\bm{m}$, is computed. Figures 4(b) and (c) display the spatial distribution of the relative error for FWI and SVGD, respectively. Within the ROI (pink dashed circle), SVGD achieves a lower maximum relative error of 0.84% and a mean error of 0.035%, compared to 1.563% and 0.064% for FWI. The results suggest that the SVGD algorithm performs well in inversion imaging, and provides an additional standard deviation map to evaluate the voxel-wise uncertainty of the SOS estimate.
Figure 5: Sensitivity of the SVGD algorithm to the number of particles. (a) The mean SOS for different particle counts ($n=3,5,10,15$, from left to right). (b) The standard deviation of the SOS.
In SVGD inversion, one important hyperparameter is the number of particles, denoted as $n$. Figure 5 (a) and (b) show the mean and the standard deviation of SOS, respectively, for different particle counts: $n=3,5,10,15$ (from left to right). It can be observed that even with a small number of particles, the mean SOS inversion using the SVGD algorithm still produces satisfactory results. As the number of particles increases, the standard deviation distribution becomes more consistent, with the boundaries between different sound speeds becoming clearer. Additionally, for higher particle counts, the standard deviation at the low-to-high speed boundaries within the tissue becomes larger. This could potentially reflect the reflection or scattering of sound waves at these boundaries, highlighting the improved precision of the inversion with more particles.
Figure 6: Uncertainty quantification of SVGD-based inversion under varying noise levels. (a) Mean SOS. (b) Standard deviation of the SOS. The inversion employs 20 SVGD particles. The left, middle, and right panels correspond to SNRs of 30 dB, 20 dB, and 10 dB, respectively. The pink dashed circle (radius = 3.27 cm) indicates the ROI.
To further assess the uncertainty quantification capability of the SVGD-based inversion method, we investigate the variation of posterior uncertainty under different levels of measurement noise. Gaussian noise with zero mean is added to the synthetic observed data, with signal-to-noise ratios (SNRs) of 30 dB, 20 dB, and 10 dB representing low, moderate, and high noise levels, respectively. For each noise level, the inversion is performed using 20 SVGD particles, initialized by adding Gaussian white noise (mean 0, standard deviation 0.01) to the initial pure water model. The prior distribution is modeled as a truncated Gaussian distribution, with $\mu_{i}=0.444~\text{s}^{2}/\text{km}^{2}$ and $\sigma_{i}=0.1~\text{s}^{2}/\text{km}^{2}$. The uncertainty is evaluated by computing the pointwise standard deviation across the particles. As shown in Figure 6(a), the mean of SOS remains relatively accurate under low and moderate noise, while performance deteriorates under high noise. The corresponding standard deviation maps in Figure 6(b) indicate increased uncertainty with decreasing SNR. The results demonstrate that SVGD captures the relationship between measurement noise and uncertainty in a consistent manner.
3.3 Breast tissue imaging using SVGD and FWI algorithms